OpenStudy (anonymous):

Find the missing elements:- l=486, r=3, Sn=728

OpenStudy (anonymous):

a=? and n=?

OpenStudy (anonymous):

l=486??what is it

OpenStudy (anonymous):

the last term

OpenStudy (anonymous):

l=ar^(n-1)

hartnn (hartnn):

2 equations, 2 unknowns...

OpenStudy (anonymous):

486=a . 3^(n-1) ------eqn (i)

hartnn (hartnn):

S = a (3^n-1)/2=728

hartnn (hartnn):

3^(n-1)=3^n /3

hartnn (hartnn):

2*728 = a (3^n)-a ---->from sum formula 3*486 =a*3^n--->from last term substitute for a*3^n and u get a from 1st equation

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

yeah i did it a=2 and n=6

hartnn (hartnn):

good :)